### v2

# Java-Class-Hijack: Software Supply Chain Attack for Java based on Maven Dependency Resolution and Java Classloading

Abstract.
We introduce Java-Class-Hijack, a novel software supply chain attack that enables an attacker to inject malicious code by crafting a class that shadows a legitimate class that is in the dependency tree. We describe the attack, provide a proof-of-concept demonstrating its feasibility, and replicate it in the German Corona-Warn-App server application. The proof-of-concept illustrates how a transitive dependency deep within the dependency tree can hijack a class from a direct dependency and entirely alter its behavior, posing a significant security risk to Java applications. The replication on the Corona-Warn-App demonstrates how compromising a small JSON validation library could result in a complete database takeover.
Software Supply Chain Attack, Java, Maven, Gradle, Namespace, Class Hijacking
copyright: none
1. Introduction
Several types of software supply chain attacks consist in executing malicious code from a dependency that is masquerading as a legitimate one. In Ladisa and colleagues’ taxonomy, these attacks fall under the category of dependency confusion (Ladisa et al., 2022). One example of this attack is the MavenGate attack (noa, 2024d). The malicious actor hijacked a legitimate dependency by purchasing an abandoned domain corresponding to the dependency, and released a new version with a backdoor. This new version was fetched by automated build systems, with no warnings, and was included in the final artifact of several projects. Such hijacking could be pervasive, as 6,170 out of 33,939 domains of dependencies in Maven Central were found to be vulnerable to this attack as of discovery of MavenGate (noa, 2024d).
In this paper, we propose a new software supply chain attack, Java-Class-Hijack, that is based on class confusion. Here the attacker crafts a malicious class that has the same name as a legitimate class in a dependency of an application. Then attacker exploits dependency resolution of Maven and how Maven packages artifacts so that the Java classloader loads the malicious class instead of the legitimate one.
We illustrate the methodology behind Java-Class-Hijack, and provide a proof-of-concept that demonstrates the feasibility of the attack. We also replicate the attack in a real-world project, cwa-server, which is the backend service of the German Coronavirus tracking application, Corona-Warn-App. Finally, we discuss that the attack is not limited to Maven and can be replicated in Gradle projects as well.
2. Background
We introduce two key steps that occur when developers build a Java project with Maven: Maven builds an artifact for the project, and the Java classloader derives the classpath from that artifact. Several design decisions made when implementing these two steps lead to the feasibility of Java-Class-Hijack, a software supply chain attack on the built project.
2.1. Maven Build Phase
The build phase of Maven operates on two inputs: the source code of the project and a configuration file that declares the dependencies of the project. This phase outputs a compiled artifact. This artifact can either be used as a dependency in other projects or can be executed as a standalone application. This phase happens in two steps: dependency resolution and packaging the artifact.
2.1.1. Dependency Resolution
This step consists of determining the exact full list of direct and transitive dependencies of the project. Each maven dependency is identified by a group ID, artifact ID, and version (G.A.V). Maven uses these three attributes to fetch the dependency from the official package registry - Maven Central - or from other repositories defined by the developer. The resolved dependencies consist of a tree where the root node is the project itself, and the children nodes are the dependencies of the project, as illustrated in Figure 1. It can happen that the same dependency (same group ID and artifact ID) appears several times in the tree but with different versions. In this case, Maven has a specific strategy to keep only one version, and avoid name conflicts later: Maven systematically chooses the dependency that appears first when going through the dependency tree in Breadth First Search (BFS) order (noa, 2024c). Once all dependencies are resolved and conflicts are resolved, Maven downloads the dependencies.
{forest}
[…]
Figure 1. Resolved dependency tree of a sample Maven project showing its direct and transitive dependencies.
2.1.2. Artifact Packaging
The next step in the build phase is to package the compiled source code of the project and all the resolved dependencies into an artifact. The artifact consists of the compiled source code of the project, together with classfiles of the dependencies or the nested JAR file of the dependency. The dependency tree is iterated in Depth First Search (DFS) Order (noa, 2023b) using Maven and all classfiles or JARs corresponding to the nodes are packaged into an artifact – which is also a JAR. For example, the DFS order of the dependency tree in Figure 1 is Project, D1, D11, D111, D112, D2, D21, D211, D22, D221. The contents of the output artifact have no standard format and will vary depending on the build plugin used. For example, Maven Shade Plugin outputs a single ‘uber jar’ (noa, 2024a) that includes all the classfiles of the project and its dependencies as shown in 1. Spring Boot Maven Plugin allows nesting the dependency JARs inside the artifact JAR (noa, 2024e) as shown in 2. However, the order of files related to the dependencies in the artifact is always in DFS order as these plugins are built on top of Maven.
1// project.jar
2+-META-INF
3| +-MANIFEST.MF
4+-Main.class
5+-D1.class // a class from D1
6+-D11.class
7+-D111.class
8+-D112.class
9+-D2.class
10+-D21.class
11+-D211.class
12+-D22.class
13+-D221.class
Listing 1: Classfiles in dependency order inside a JAR file of the sample Maven project packaged using Maven Shade Plugin.
1// project.jar
2+-BOOT-INF
3 +-classes
4 | +-Main.class
5+-BOOT-INF
6 +-lib
7 | +-D1.jar
8 | +-D11.jar
9 | +-D111.jar
10 | +-D112.jar
11 | +-D2.jar
12 | +-D21.jar
13 | +-D211.jar
14 | +-D22.jar
15 | +-D221.jar
Listing 2: JAR file of dependencies order inside a JAR file of the sample Maven project packaged using Spring Boot Maven Plugin.
2.2. Java Classpath
Java classpath is a string of locations of Java classes and/or JAR files where code can be found to run a Java application. It is dependent upon how the artifact is built and how the classes are ordered inside the artifact. In the following sections, we explain how Java classpath is built and how Java classloaders use it to load classes at runtime.
2.2.1. Building Classpath
Java classpath is a string concatenating the locations of compiled Java code (or classfiles) and JAR files. The purpose of the classpath is to provide the Java runtime environment with the location of the classes required to run the application. These classes come from three sources: classes written by the developer, classes that are part of declared dependency, and classes from the Java standard library (Sharma et al., 2024). Since the classes from the Java standard library are always present in the Java runtime environment, the classpath string only needs to contain the locations of the developer-written and dependency classes.
To make the classpath available to the Java runtime environment, Java provides multiple ways to set the classpath. One of them that we will be using in this paper is the -jar flag. The -jar flag is used to execute the Java application from a JAR file. For example, to execute the application in Figure 1, we would need to run java -jar project.jar. This would do two things: 1) execute Main.class as it is the entry point defined in MANIFEST.MF and 2) set the classpath string to all the dependency classes inside the JAR. The classpath string in this case would look like: "D1.class; D11.class; D111.class; D112.class; D2.class; D21.class; D211.class; D22.class; D221.class" or "D1.jar; D11.jar; D111.jar; D112.jar; D2.jar; D21.jar; D211.jar; D22.jar; D221.jar" depending upon how project.jar is packaged. Its order is the same as how classes/JAR files were ordered in the artifact. Finally, once the application is running, the Java classloader will use the classpath to look up the classes required at runtime for execution.
2.2.2. Class Lookup in Classpath
Java classloaders are responsible for looking up and loading classes at runtime. By default, three built-in classloaders are used to load classes: Bootstrap Classloader, Extension Classloader, and System Classloader (noa, 2023a). We will focus on the System Classloader as it is the one that is used to load classes other than those belonging to Java standard library.
System Classloader performs a linear search on the classpath for the fully qualified name of the class (noa, 2021, §"Finding the Right Class") (Wang et al., 2022) by the classloader. For example, to look up the class D2.class, the classloader will read the project.jar file in the classpath. Depending upon how the JAR file (1 or 2) is packaged, the classloader will either read the D2.class file directly from the JAR file or read the D2.jar file from the JAR file and then read the D2.class file from the D2.jar file. If there are two occurrences of the same class, the first that appears in the path is selected for further execution. It is clear that being first on the classpath allows one to shadow any class implementation that is included afterward.
3. Attack Description
Java-Class-Hijack exploits the specifics of how dependencies are packaged into an artifact by Maven and then how the classpath is looked up for classes. The attack consists of a malicious class masquerading a legitimate class, compromising the integrity and security of the resulting software artifact. This is possible because Maven does not enforce correspondence between Maven artifact names and the names of classes they include. This implies that anyone can publish a class under a popular package name inside a random Maven artifact (noa, 2024b).
The attack takes place in two steps. The first step for this attack is to craft a malicious class that has the same fully qualified name as a legitimate class that is going to be packaged in the built project. The second step is to embed this malicious class in a dependency that will be included in the classpath before the dependency that includes the legitimate class. Recall that, if there are multiple classes with the same fully qualified name in different dependencies, the class that appears first in the DFS order is the one that is executed at runtime. The final JAR file includes the classes of all resolved dependencies.
{forest}
[…]
Figure 2. Infected dependency tree of a sample Maven project showing Genuine and Malicious classes.
Figure 2 shows an example of the attack described above. D111 is a transitive dependency containing MaliciousClass, while D2 is a direct dependency containing GenuineClass. During DFS iteration, MaliciousClass in D111 is found first and packed into the final JAR file before GenuineClass in D2. Consequently, at runtime, the malicious class is loaded from the classpath instead of the genuine class. More generally, all dependencies to the left of the target dependency, D2 in this case, can be attacked by ‘hijacking’ the genuine class name of one of the classes in the target dependency. Popular classes, for example org.apache.commons.lang3.StringUtils, are more likely to be targeted by this attack, as they are more likely to be included in the final artifact. Such an attack is stealthy since the malicious class can be located deep in the dependency tree which would make it hard to discover.
4. Ways to Hijack
Java-Class-Hijack consists in crafting a malicious class and then exploit Maven and Java design, in order to have the malicious class replace a legitimate one. In this section, we describe two strategies that can lead to the malicious class be incorporated in the dependency tree of an application.
Creation of Malicious Dependency: The attacker develops a malicious library with classes that have the same fully qualified name as those of the genuine dependency, and uploads it to the central Maven repository The purpose of the malicious dependency can be attractive to users which encourages its integration into client projects. Another option for the attacker is to create a package with a name similar to an existing one, leading to a typosquatting attack (Ohm et al., 2020). Typosquatting exploits user errors in typing package names, causing them to inadvertently download and use the malicious package instead of the legitimate one. If a client declares the malicious dependency in the "correct position" in the dependency tree, Maven prioritizes it in the depth-first search (DFS) order, including the malicious classes in the final artifact instead of the legitimate ones.
Library takeover: The attacker takes control of a dependency as a maintainer. Additionally, the attacker can take control of the dependency domain in cases where the dependency has not been maintained for a long time  (noa, 2024d). Having become a maintainer, the attacker introduces compromised versions of the library. Users rely on the integrity of the original library and, by automatically updating, download and use the malicious versions in the client application. This is effective in less popular libraries, which tend to have less oversight.
5. Replication
We present two proofs-of-concept: the first is entirely constructed from scratch to effectively illustrate the phases of the class hijacking attack. The second is based on a real, open-source project available on GitHub repository called Corona-Warn-App (cwa, 2023) and is used to highlight the potential impact of the attack.
All the proof-of-concepts, attack replication and experiments with other languages are available on our GitHub repository (poc, 2024).
5.1. Abstract proof-of-concept
In the first proof-of-concept, we have a victim application that has two dependencies: org.test.nicelibrary and org.evil.attackerlibrary. Simultaneously, org.evil.attackerlibrary depends on another artifact named org.evil.fakelibrary, which includes a malicious implementation of org.test.NiceClass, also originally defined in org.test.nicelibrary.
1// dependencies of the victim project
2+-org.test.nicelibrary
3| +-org.test.NiceClass // Genuine
4+-org.evil.attackerlibrary
5| +-fakelibrary
6| | +-org.test.NiceClass // Malicious
Listing 3: Locations of org.test.NiceClass
As explained in the previous sections, Java determines which implementation to use by examining the classpath, which Maven defines using a depth-first approach. Therefore, the order of dependencies in the pom.xml file can result in different outputs after compiling and packaging the victim application.
If org.test.nicelibrary is defined before the attacker library (org.evil.attackerlibrary), like in Figure 3, the genuine implementation of org.test.NiceClass will be used by the victim, and the attack will not succeed.
1<dependency>
2 <groupId>org.test</groupId>
3 <artifactId>
4 nicelibrary
5 </artifactId>
6 <version>1.2</version>
7</dependency>
8
9<dependency>
10 <groupId>org.evil</groupId>
11 <artifactId>
12 attackerlibrary
13 </artifactId>
14 <version>1.0</version>
15</dependency>
(a) Dependencies defined in the pom file: attackerlibrary is after nicelibrary
{forest}
[…]
(b) Tree representation of the resolved dependencies
Figure 3. Victim dependency tree for a not successful attack: NiceClass is loaded from the nicelibrary
However, if org.evil.attackerlibrary is defined before org.test.nicelibrary in the pom.xml file, like in Figure 4, then the malicious implementation of org.test.NiceClass from the transitive dependency org.evil.fakelibrary will be used. This allows a transitive dependency to hijack a class from another dependency defined on the right side of the tree. This malicious class can modify the overall application behaviour, steal data or include any kind of vulnerability, without the need to attack the specific dependency that defines the legitimate class. It is important to notice that the malicious dependency can be at any depth level in the dependency tree.
1<dependency>
2 <groupId>org.evil</groupId>
3 <artifactId>
4 attackerlibrary
5 </artifactId>
6 <version>1.0</version>
7</dependency>
8
9<dependency>
10 <groupId>org.test</groupId>
11 <artifactId>
12 nicelibrary
13 </artifactId>
14 <version>1.2</version>
15</dependency>
(a) Dependencies defined in the pom file: attackerlibrary is before nicelibrary
{forest}
[…]
(b) Tree representation of the resolved dependencies
Figure 4. Victim dependency tree for a successful attack: NiceClass is loaded from the fakelibrary
5.2. Replication in cwa-server
To assess the potential attack impact on a real project, we replicated the attack on the Corona-Warn-App backend services, referred to as cwa-server (cwa, 2023). Corona-Warn-App is Germany’s Coronavirus tracking application, used during the 2020 pandemic to perform contact tracing using iOS and Android features. The backend services are written using the popular Java framework Spring-Boot and, since Corona-Warn-App is no longer in use, the repository is currently archived. This makes it an ideal base for a proof-of-concept that would also apply to other Java web applications.
{forest}
[…]
Figure 5. Truncated dependency tree for Corona-Warn-App backend service, from the parent pom.xml file
Corona-Warn-App is composed of many services, each with specific dependencies, but all depending on a parent pom.xml file with some common dependencies. Being able to compromise the parent dependency tree means that any service will also be vulnerable to the attack.
In the parent dependency tree, a small (i.e. < 1K stars) dependency (everit-json-schema) is defined before the database SDK (postgresql). Consequently, an attacker who can include a malicious class in everit-json-schema, or in any of its dependencies, can control the implementation for any class from the postgresql SDK, effectively overtaking the database connection layer.
In this proof-of-concept, we modified the parent pom.xml file to point to a compromised version of everit-json-schema, simulating a library takeover. In this compromised version, we made only one change: we added one malicious dependency. This dependency defines a modified implementation of the org.postgresql.Driver class. Our implementation simply throws an exception to facilitate an easy check for a successful exploit.
When the service starts, Spring-Boot loads the database connection class (org.postgresql.Driver) from the classpath. Since the malicious version of this class is found on the left side of the dependency tree, it is on the classpath and it is loaded at runtime, which leads to the exception being thrown.
6. Class Hijacking in Gradle ecosystem
Java-Class-Hijack exploits specific design decisions in Maven related to packaging artifacts, classpath resolution, and the default Java class look up algorithm. In this section, we discuss the opportunity for the same attack on an application that builds with Gradle, an alternative build system for Java.
Gradle made different design design decisions compared to Maven, regarding the construction of the dependency tree and of the final classpath. First, the classpath is generated using a breadth-first search algorithm. This means that the direct dependencies are included first. This reduces the attack surface as the attacker library can only hijack classes from libraries at the same level or below, always on the right side of the tree. Second, additional repositories for dependencies are ignored by Gradle, requiring manual insertion. This prevents the attacker from hiding malicious code in a self-managed repository, which might otherwise evade extra checks.
The last step of the attack assumes a linear search in the classpath when loading a class. This depends on Java and is independent of Gradle or Maven.
In summary, Java-Class-Hijack  is feasible on an application that builds with Gradle. Yet, it is significantly more challenging than with Maven, due to different key design decisions.
7. Mitigation
Even though the industry aims to avoid compromised dependencies in the tree at all, limiting the potential impact of any such dependency on the project should also be an important goal for software engineers. This section presents several possible mitigations for this attack.
7.1. Maven Enforcer Plugin
In Java projects built with Maven, developers could use the maven-enforcer-plugin (mav, 2024) with the banDuplicateClasses (ban, 2024) extra rule enabled. This plugin enforces specific rules at build time and adds extra checks. The banDuplicateClasses rule, in particular, fails the build process when a class collision is detected. However, this rule may increase build time and network usage and can result in false positives, as class collisions are sometimes expected and not indicative of an attack. For example, in the most recent versions of hamcrest, its packages have been published in separated artifacts, while before they were deployed into a single one. When the new version is included, a class collision may happen as Maven dependency resolution is based on the artifact id and not on package id.
7.2. Sealed JARs
A project maintainer can enhance the security of dependent applications by creating a sealed JAR (sea, 2024) for the packages defined by the project itself. When an application uses a sealed JAR and the same package name is redefined, a runtime exception is thrown, thereby reducing the impact of Java-Class-Hijack on the application. However, this is not sufficient, as a complete copy of the original dependency could be included in the malicious dependency, bypassing the sealed JAR definition checks. For example, if the victim only calls org.nice.ClassA, the attacker needs to create a complete new malicious dependency with org.nice.ClassA along with all the classes that are called by org.nice.ClassA in the benign package. If org.nice.ClassA calls org.nice.ClassB, and if the attacker does not include ClassB in the dependency, it will result in a RuntimeException.
7.3. Java Modules
With Java 9, Project Jigsaw (jav, 2024) was introduced into the JDK, adding the concept of modularity to the Java ecosystem. Java Modules facilitate the maintenance of large projects and libraries while enhancing the security of the Java SE Platform. A modularized application is immune to Java-Class-Hijack because a package collision results in an automatic compilation failure. Since defining a module for an application only requires adding a module-info.java file in the project root, this method is recommended to increase security and prevent Java-Class-Hijack. To mitigate the attack, is sufficient that the victim project is defined as a module, the dependencies can also be modules but is not required.
In 4 an example of module definition for our abstract PoC is presented, this results in a compilation failure as org.test.NiceClass is defined in both fakelibrary and nicelibrary.
1module victim {
2 uses org.test.NiceClass;
3 requires org.evil.attackerlibrary;
4 requires fakelibrary;
5 requires nicelibrary;
6}
Listing 4: Example of module-info.java to mitigate the attack
8. Conclusion & Takeaways
We introduced a novel software supply chain attack, Java-Class-Hijack, that targets Java projects built with Maven. We provided a proof-of-concept that illustrates the attack and demonstrated the feasibility of the attack by replicating it in a real-world project, cwa-server.
The key takeaways of this paper are:
  • Group ID and Artifact ID do not constrain the Java package IDs that can be included in the artifact. Consequently, different Maven artifacts could contain two instances of the same class, one of which could be malicious.
  • Default Java classloader picks the first class that it finds in the classpath and is blind toward possible naming collisions. Consequently, if attackers succeed in inserting a malicious class before the legitimate version on the classpath, they can alter the application’s behavior.
  • Java-Class-Hijack is stealthy as the malicious class can be buried deep in the dependency tree so it is hard to determine the correct declaration order.
  • Java Modules, when correctly defined, block this type of attack.
  • Projects built with other ecosystem are also vulnerable to Java-Class-Hijack.
Acknowledgements.
We thank Hervé Boutemy, who is a PMC member of The Apache Software Foundation, for giving ideas about mitigations. This work was supported by the CHAINS project funded by the Swedish Foundation for Strategic Research (SSF), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) and the IVADO consortium.